This problem in extra work but i am confused. If anyone can help thanks.
Solve algebraically, using the five step method. Show all steps.
The width of a poool is 3/4 of its length. It takes 126 feet of fence to edge the pool What are the dimensions of the pool?

Answer by stanbon(75887)

The width of a poool is 3/4 of its length. It takes 126 feet of fence to edge the pool What are the dimensions of the pool?
-------------------------
Let the length be L
Then the width = (3/4)L
------------------------------
Perimeter = 2(length + width)
126 = 2(L + (3/4)L)
63 = (1 3/4)L
63 = (7/4)L
L = (4/7)63
L = 4*9
L = 36 feet (length of the pool)
W = (3/4)L = (3/4)36 = 27 feet (width of the pool)
======================================================
Cheers,
Stan H.